What number of permutation of the letters in the word smart?
For the first letter, you can have any of 5 choices. Then for the next, you can have any of four. Then for the next, three and so on. Thus the number of permutations can be calculated by 5x4x3x2x1. Doing this gives 120.
Therefore the number of permutations of the letters in the word smart is 120.
Can probability outcomes be expressed as negative integer's?
The outcomes of a trial can be a negative integer but the probability of any outcome must be in the range [0, 1].
What is the probability of getting at least two tails?
That depends how many times you flip the coin.
What denotes the number of standard deviations a particular score is from the mean?
The absolute value of the z-score.
How many ways can you arrange 10 trophies on a table in a rectangular array?
There are four ways to arrange 10 objects in a rectangular array; 1 x 10, 2 x 5, 5 x 2, and 10 x 1. Two of these patterns are mirror images of two others.
What is the probability of flipping 3 coins at the same time and getting all heads or all tails?
When you flip three fair coins, there are eight possible ways -- or states -- they can attain. Because we have only three coins, it's easy to list all the states. Here they are:
T T T
T T H
T H T
T H H
H T T
H T H
H H T
H H H
Each of those states has an equal chance of occurring. So the TTT state has a one-in-eight chance. That can be expressed mathematically as 1/8 = 0.125. The HHH state has the same chance: 0.125.
Since you asked for the probability of getting all heads OR all tails, you would add the two probabilities together: 0.125 + 0.125 = 0.25. In other words, there is a one-in-four (1/4) chance of flipping all heads OR all tails. (Remember that 1/8 + 1/8 = 1/4 = 0.25)
How many ways can 24 objects be arranged?
They can be arranged in the following patterns...
OOOOOOOOOOOOOOOOOOOOOOOO
OOOOOOOOOOOO
OOOOOOOOOOOO
OOOOOOOO
OOOOOOOO
OOOOOOOO
OOOOOO
OOOOOO
OOOOOO
OOOOOO
OOOO
OOOO
OOOO
OOOO
OOOO
OOOO
OOO
OOO
OOO
OOO
OOO
OOO
OOO
OOO
OO
OO
OO
OO
OO
OO
OO
OO
OO
OO
OO
OO
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
O
What are the odds of drawing a spade from a deck of 52 cards?
For a dingle card drawn at random, it is 1/4.
What is a probability of getting a sum Nine when two dice are thrown?
Throwing two dice, there are a total of 36 combinations (1-1, 1-2, 1-3, etc.), each of which has the same probability. Of these, the following give a sum of nine: 3-6, 4-5, 5-4, 6-3. That makes four favorable cases, or a probability of 4/36.
When rolling five dice what is the probability that all 5 dice will be the same number?
The probability of rolling the same number on five dice is (1/6)4, or about 0.0007716.
What is the probability of randomly selecting a three-children family with exactly two girs?
It is 3/8.
What is the probability of randomly selecting a three-child family with exactly two girl children?
2/6 is not accurate. using a theoretical method for equally likely outcomes, there are 2 possible outcomes for each birth: either a boy(B), or a girl (G). For a family of three children, the total number of possibilities (birth orders) is 2*2*2=8 to double check this work, here are the eight possible outcomes:
BBB, BBG, BGG, GBB, GBG, GGB, and GGG. You want EXACTLY two girls, this assumes that the other must be a boy. Therefore, the probability that a three child family has 2 girls one boy is P(2 girls)=3/8=0.375
What is the volume of a six sided dice?
it all depends on the length the width and the height. if you know those then just add them all up and you will have the volume.
Using the hypergeometric distribution, the answer is 2114/3003 = 0.7040
What is probability that two heterozygous parents would have an offspring that produced round seeds?
It is a 75% chance that the seeds will be round.
How can you find a probability expressed as a percentage from a punnett square?
You can find probability form a Punnett square by turning fractions into percents
How many ways can you arrange 7CDs on a shelf?
The number of permutations of 7 things taken 7 at a time is 7 factorial, or 5040.
How do you convert probabilities to odds?
Just take the probability, which is a decimal number between 0 and 1, and convert it into a fraction. For instance, a probability of 0.75 corresponds to odds of 3 in 4.
What is the probabilility that the spinner will land on six?
The probability that the spinner will land on six depends on how many numbers are on the spinner. If the spinner is only 1 through 6, then there is a 16.67% probability that the spinner will land on six with each spin.
It is 6!/3! = 6*5*4 = 120 ways